OpenStudy (anonymous):

g(x) = 4x - x^2 Find the equations of the two tangent lines to the graph of g that pass through P(2,5), which is not on the graph g(x)

OpenStudy (anonymous):

draw a graph of g(x) should g not be g(x) = 4x² − x or something, because otherwise the question makes no sence (g would be a straight line)

OpenStudy (anonymous):

Nice catch. Edited.

OpenStudy (anonymous):

Hang on, I'm in the middle of answering another question

OpenStudy (anonymous):

Ok, we have g(x) = 4x - x^2 Find the equations of the two tangent lines to the graph of g that pass through P(2,5), which is not on the graph g(x) Just a sec, need a spreadsheet to create a nice graph hang on

OpenStudy (anonymous):

Can I do this without a graph?

OpenStudy (anonymous):

Preferably with the first derivative.

OpenStudy (anonymous):

Yest that's right

OpenStudy (anonymous):

The tangent lines have the direction of the derivative of g at some point on g

OpenStudy (anonymous):

Here is what I got the tangents have a direction of dg/dx and are straight lines y = ax + b where a is dg(x)/dx | x on g and x on the tangent line as well g'(x) = 4 - 2x_g, x_g is an x on the parabola x so y_t = (4 - 2x_g)x_t + b the tangent line passes through (2, 5) 5 = (4 - 2x_g)2 + b 5 = 8 -4x_g + b b = 4x_g - 3 There is also a point on the tangent line and the parabola: (x_t, y_t) = (x_g , y_g), y_t = (4 - 2x_g)x_t + b y_g = 4x_g - x²_g x_g = x_t and y_t = y_g, so (4 - 2x_g)x_g + b = 4x_g - x²_g 4x_g -2x²_g +b = 4x_g - x²_g b = x²_g We already had an expression for b so 4x_g - 3 = x²_g x²_g - 4x_g +3 = 0 so x is (4 + sqrt(16 - 12))/2 = 3 and (4 - sqrt(16 - 12))/2 = 1 so the first tangent line is y = (4 - 2x_g)x + x²_g , x_g = 3 y = (4 - 2*3)x + 9 y = -2x + 9 and the second tangent line is y = (4 - 2x_g)x + x²_g , x_g = 1 y = 2x + 1

OpenStudy (anonymous):

Yes you can do it without a graph, but a graph shows what is going on and shows the anwers make sense A sketch in advance is also handy  I usesed this site to make the graphs http://rechneronline.de/function-graphs/